# Caching-Augmented Lifelong Multi-Agent Path Finding

### v2

@@ V Experimental Results
We use Lifelong MAPF without cache as a baseline, which aligns with our problem definition. To evaluate performance, we compare CAL-MAPF with Lifelong MAPF, using LaCAM as the low-level MAPF solver for generating collision-free paths. Both CAL-MAPF and Lifelong MAPF were implemented in C++, building on parts of the existing LaCAM codebase [^1] The LaCAM source code is publicly accessible at https://github.com/Kei18/lacam. Our implementation is available at https://github.com/HarukiMoriarty/CAL-MAPF. All experiments were conducted on a system running Ubuntu 20.04.1, equipped with an AMD Ryzen 3990X 64-core CPU and 64GB RAM at 2133 MHz.
@@ V-A Test Settings
As shown in Figure 1, we designed a warehouse map (27x71) with caches based on our problem definition. The map is adapted from the warehouse map of MAPF benchmark [2]. It includes 1600 shelf grids, a maximum of 80 cache grids and 4 unloading ports. We tested CAL-MAPF in both multi-port and single-port scenarios, as depicted in Figure 1. The multi-port scenario has 4 working groups of unloading ports and cache grids, each with a maximum of 20 cache grids. The single-port scenario maintains the same number of total cache grids and agents as the multi-port but only has one unloading port. Each shelf grid represents a unique kind of item, and we randomly assign an index to each shelf grid. The maximum cache-to-shelf ratio is 5%. We test all scenarios with different cache numbers {16, 32, 48, 64, 80} by deleting some cache grids in Figure 1. We have also chosen various total numbers of agents, $\{4,8,16,32,64,128,192,256\}$. In the multi-port scenario, each group is assigned an equal number of agents $\{1,2,4,8,16,32,48,64\}$.
Since we can expect the distribution of the task queue could significantly affect the performance of cache design, we designed three input task distributions to test CAL-MAPF, including: (1) $M$-$K$ distribution (MK): For any consecutive subarray of length $M$ in the task queue $Q$, there are at most $K$ different kinds of item. This distribution is inspired by [37], where LRU has been proven to have a limit on the cache miss rate and to be better than FIFO. This distribution can also represent that there are only several items people purchase daily, and some items may become very popular at one time, replacing previously popular items. (2) 7:2:1 distribution (Zhang): there are 70% kinds of items with only a 10% appearance probability in the task queue, 20% kinds of items with a 20% probability, and 10% with a 70% appearance probability [48]. (3) Real Data Distribution (RDD): we obtain data from Kaggle’s public warehouse222kaggle.com/datasets/felixzhao/productdemandforecasting, build a probability distribution based on the frequency of Kaggle data items and generate tasks from this probability distribution.
@@
We randomly generate agent start locations and $Q$ for all distributions with a total length of 1,000. For the MK distribution, we select $M=200$ and $K$ values of $\{1,20,40,80,120,160\}$. We allocate a total of 10 seconds for the low-level MAPF solver to find a collision-free solution for all 1,000 tasks in $Q$. We show performance using makespan which is the completion time when the last task has been accomplished. For CAL-MAPF, we use three different cache replacement policies. Least Recently Used (LRU), First-In-First-Out (FIFO), and RANDOM. We use NONE to represent the baseline method.
@@ V-B Performance
Many variables can influence the performance of CAL-MAPF, including the number of agents, the number of caches, and the distributions of input tasks. As illustrated in Figure 4, increasing the number of caches leads to improvements in cache hit rate and makespan performance for both single- and multi-port scenarios. And the cache hit rate decreases as the number of agents grows. It is also observed that as the number of agents continues to increase (from 64 to 256), there is not much difference in the makespan for both CAL-MAPF and the baseline, and the cache hit rate is dropping.
Two reasons could contribute to this scenario: (1) CAL-MAPF and baseline’s makespan remain at the same level or even worse with too many agents: traffic congestions occur in the map as the number of agents increases. (2) The CAL-MAPF cache hit rate continues to decrease as the makespan performance shows minimal differences: We employ a lock mechanism to maintain items stored in the cache area. Traffic congestion prevents many agents from reaching their target in the cache and agents always hold locks until they reach their target locations. For other agents not significantly affected by traffic congestion, they are constantly assigned new tasks and cannot obtain locks in the cache. The item distribution cannot update with input tasks, which ultimately leads to a continuous drop in the cache hit rate. This second reason can also clarify why CAL-MAPF performs better as the number of caches increases: enlarging the cache area mitigates traffic congestion and enhances the cache hit rate.
@@
Furthermore, in Figure 5, an abnormal increase in makespan is observed at baseline under the Zhang distribution with 256 agents. Traffic congestion could contributing to this abnormal poor performance. Figure 3 illustrates the average frequency of agent wait actions on each grid when the number of agents reaches 256 with the Zhang distribution. It is evident that both CAL-MAPF and baseline experience severe traffic congestion.
High cache hit rates and smooth traffic are crucial for the performance of CAL-MAPF. To enhance the cache hit rate, we can increase the number of caches. And we can also directly add more agents to improve the makespan performance. However, both methods come with their drawbacks. The number of caches is constrained by the space available close to the unloading port. Introducing more agents can lead to severe traffic congestion, ultimately causing the hit rate to drop. Nevertheless, there are potential solutions to mitigate these adverse effects, such as implementing more efficient map designs [13] and using one-way systems near the cache and unloading ports [14]. Additionally, as we currently employ a simple task assignment policy, we could implement a more advanced policy with predictive capabilities, leveraging real warehouse task data to enhance the cache hit rate. The cache replacement policies, such as LRU and FIFO, may be overly simplistic for CAL-MAPF. Incorporating more complex policies, including learning-based approaches, could further improve the cache hit rate, especially since CAL-MAPF allows more planning time than traditional caches in computer architecture.
@@ VI Conclusion
This work presents a novel mechanism, named Caching-Augmented Lifelong MAPF (CAL-MAPF), designed to improve the performance of Lifelong MAPF. We have developed a new map grid type called cache for temporary item storage and replacement. Additionally, we have devised a locking mechanism for caches to improve the stability of the planning solution. This cache mechanism was evaluated using various cache replacement policies and a spectrum of input task distributions. Through experimentation, we identified three main factors that significantly impact CAL-MAPF performance: suitable input task distribution, high cache hit rate, and smooth traffic. In general, CAL-MAPF has demonstrated potential for performance improvements in certain task distributions and map and agent configurations.

### v3

@@ V Experimental Results
We use Lifelong MAPF without cache as a baseline, which aligns with our problem definition. To evaluate performance, we compare CAL-MAPF with Lifelong MAPF, both using LaCAM as the MAPF solver for generating collision-free paths. CAL-MAPF and Lifelong MAPF were implemented in C++, building on parts of the existing LaCAM codebase [^1] The LaCAM source code is publicly accessible at https://github.com/Kei18/lacam. Our implementation is available at https://github.com/HarukiMoriarty/CAL-MAPF. All experiments were conducted on a system running Ubuntu 20.04.1, equipped with an AMD Ryzen 3990X 64-core CPU and 64GB RAM at 2133 MHz.
@@ V-A Test Settings
As shown in Figure 1, we designed a warehouse map (27x71) with caches based on our problem definition. The map is adapted from the warehouse map of MAPF benchmark [2]. It includes 1600 shelf grids, a maximum of 80 cache grids and 4 unloading ports. The maximum cache-to-shelf ratio is 5%. We tested CAL-MAPF in both multi-port and single-port scenarios, as depicted in Figure 1. The multi-port scenario has 4 working groups of unloading ports and cache grids, each with a maximum of 20 cache grids. The single-port scenario maintains the same number of total cache grids and agents as the multi-port but only has one unloading port. Because each shelf grid represents a unique kind of item, and we randomly assign an index to each shelf grid. We test all scenarios with different cache numbers {16, 32, 48, 64, 80} by deleting some cache grids (described in Figure 1). We have also chosen various total numbers of agents, $\{4,8,16,32,64,128,192,256\}$. In the multi-port scenario, each group has an equal number of agents $\{1,2,4,8,16,32,48,64\}$.
Since we can expect the distribution of the task queue could significantly affect the performance of cache design, we designed three input task distributions to test CAL-MAPF, including: (1) $M$-$K$ distribution (MK): For any consecutive subarray of length $M$ in the task queue $Q$, there are at most $K$ different kinds of items. This distribution is inspired by [37], where LRU has been proven to have an upper bound on the cache miss rate and to be better than FIFO. This distribution can also represent several items people purchase daily, and some items may become very popular at one time, replacing previously popular items. (2) 7:2:1 distribution (Zhang): There are 70% kinds of items with only a 10% appearance probability in the task queue, 20% kinds of items with a 20% probability, and 10% with a 70% appearance probability [48]. (3) Real Data Distribution (RDD): We obtain data from Kaggle’s public warehouse data222kaggle.com/datasets/felixzhao/productdemandforecasting, build a probability distribution based on the frequency of data and generate tasks from this probability distribution.
@@
We randomly generate agent start locations and $Q$ for all input task distributions with a total length of 1,000. For the MK distribution, we select $M=200$ and $K$ values of $\{1,20,40,80,120,160\}$. We allocate a total of 10 seconds for the MAPF solver to find a collision-free solution for all 1,000 tasks in $Q$. We use makespan, the completion time when the last task has been accomplished, to show performance. For CAL-MAPF, we have tested three different cache replacement policies. Least Recently Used (LRU), First-In-First-Out (FIFO), and RANDOM. We use NONE to represent the baseline method.
@@ V-B Performance
Many variables can influence the performance of CAL-MAPF, including the number of agents, the number of caches, and the distributions of input tasks. As illustrated in Figure 4, increasing the number of caches improves cache hit rate and makespan performance for both single- and multi-port scenarios. And the cache hit rate decreases as the number of agents grows. It is also observed that as the number of agents continues to increase (from 64 to 256), there is not much difference in the makespan for both CAL-MAPF and the baseline, and the cache hit rate is dropping.
Two reasons could contribute to this scenario: (1) CAL-MAPF and baseline’s makespan remain at the same level or even worse with too many agents: traffic congestions occur in the map as the number of agents increases. (2) The CAL-MAPF cache hit rate continues to decrease as the makespan performance shows minimal differences: We employ a lock mechanism to maintain items stored in the cache area. Traffic congestion prevents many agents from reaching their target in the cache. For other agents not significantly affected by traffic congestion, they are constantly assigned new tasks. They cannot obtain locks in the cache because locks are held by agents in congestion. So the item distribution cannot update with input tasks, which ultimately leads to a continuous drop in the cache hit rate. This second reason can also clarify why CAL-MAPF performs better as the number of caches increases: enlarging the cache area mitigates traffic congestion and enhances the cache hit rate.
@@
Furthermore, in Figure 5, an abnormal increase in makespan is observed at baseline under the Zhang distribution with 256 agents. Traffic congestion could contribute to this abnormally poor performance. Figure 3 illustrates the average frequency of agent wait actions on each grid when the number of agents reaches 256 with the Zhang distribution. Both CAL-MAPF and baseline experience severe traffic congestion.
High cache hit rates and smooth traffic are crucial for the performance of CAL-MAPF. We can increase the number of caches to enhance the cache hit rate. And we can also directly add more agents to improve the makespan performance. However, both methods come with their drawbacks. The number of caches is constrained by the space available close to the unloading port. Introducing more agents can lead to severe traffic congestion, ultimately causing the hit rate to drop. Nevertheless, there are potential solutions to mitigate these adverse effects, such as implementing more efficient map designs [13] and using one-way systems near the cache and unloading ports [14]. Additionally, as we currently employ a simple task assignment policy, we could implement a more advanced policy with predictive capabilities, leveraging real warehouse task data to enhance the cache hit rate. The cache replacement policies, such as LRU and FIFO, may be overly simplistic for CAL-MAPF. Incorporating more complex policies, including learning-based approaches, could further improve the cache hit rate, especially since CAL-MAPF allows more planning time than traditional caches in computer architecture.
@@ VI Conclusion
This work presents a novel mechanism, Caching-Augmented Lifelong MAPF (CAL-MAPF), designed to improve the performance of Lifelong MAPF. We have developed a new map grid type called cache for temporary item storage and replacement. Additionally, we have devised a locking mechanism for caches to improve the stability of the planning solution. This cache mechanism was evaluated using various cache replacement policies and a spectrum of input task distributions. We identified three main factors that significantly impact CAL-MAPF performance through experimentation: suitable input task distribution, high cache hit rate, and smooth traffic. In general, CAL-MAPF has demonstrated potential for performance improvements in certain task distributions and map and agent configurations.